OpenStudy (anonymous):

Solve by substitution. 4x − 12y = 5 x = 3y + 1 (1, 4) (0, 1) No solution (negative-1, negative-2) @iGreen

OpenStudy (anonymous):

Question : 4x − 12y = 5 x = 3y + 1 Solving for y : 4(3y + 1) - 12y = 5 12y + 4 - 12y = 5 12y - 12y = 1 y = 1 Solving for x : 3(1) + 1 3 + 1 x = 4 ANSWER : (1,4)

OpenStudy (anonymous):

Solve by substitution. 4x − y = 12 negative-2x + 2y = 6 (1, 4) (5, 8) (2, 2) (negative-1, negative-3) @Kaydex

OpenStudy (anonymous):

no, its just -1,-3

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@iGreen